1. Hannah invested Php 75,000 for a pastry business. This amount is 25% of her total savings. How much is hannah's total savings?
2. TripleK Cafè has 360 regular customers. If 75% of its customer are female, how many male customers does it have
3. As a part time employee, Michelle earns Php 85 per hour. Jons earns 30% more than what Michelle earns. How much does jona earn in an hour?
4. Mang Juan has 2 baskets of mangoes to sell. If 45% of the mangoesare unripe and there are 20 mangoes in all, how many ripe mangoes are there?
5. You have a buy and sell business. You bought a second hand car for Php 120,000. You sold it to a friend at 5% higher than the price you paid for it
How much did your friend pay you for the car?
1
Expert's answer
2021-09-30T14:16:57-0400

1.

If 75000 "="25 "\%"

?"=" 100"\%"

"\frac{100\times75000}{25}"

Php "=" 300, 000

2.

360 regular customers

75% are female

So 100% = 360

25%"=" ?

"\frac{25\times360}{100}"

Answer is 90 male customers.

3.

Michelle earns 85 per hour

Moms earns 30% more than Michelle

So;

85"=" 100%

?"=" 130%

"\frac{130\times85}{100}"

Answer is 110.50

4.

100%"=" 20

55%"=" ?

"\frac{55\times20}{100}"

Answer is 11 ripe mangoes.

5.

120, 000"=" 100%

?"=" 105%

"\frac{105\times120000}{100}"

Answer is Php 126, 000